y/y-2 + y/y^2-4 = y+3/y+2
Answer by joey_quizan(15)

Given:
y/(y-2) + y/(y^2-4) = y + 3 /(y+2)
Solution:
y/(y-2) + y/[(y-2)(y+2)] = (y+3)/(y-2)
(y-2)(y+2){y/(y-2) + y/[(y-2)(y+2)] = (y+3)/(y-2)}
y(y+2) + y = (y+3)(y+2)
y^2 + 2y + y = y^2 + 5y + 6
y^2 - y^2 + 2y + y - 5y - 6 = 0
-2y - 6 = 0
y + 3 = 0
y = -3
answere: y = -3